# The Illusion of Diminishing Returns: Measuring Long Horizon Execution in LLMs

### v1

@@ Abstract
Does continued scaling of large language models (LLMs) yield diminishing returns? Real-world value often stems from the length of task an agent can complete. We start this work by observing the simple but counterintuitive fact that marginal gains in single-step accuracy can compound into exponential improvements in the length of a task a model can successfully complete. Then, we argue that failures of LLMs when simple tasks are made longer arise from mistakes in execution, rather than an inability to reason. We propose isolating execution capability, by explicitly providing the knowledge and plan needed to solve a long-horizon task. We find that larger models can correctly execute significantly more turns even when small models have 100% single-turn accuracy. We observe that the per-step accuracy of models degrades as the number of steps increases. This is not just due to long-context limitations—curiously, we observe a self-conditioning effect—models become more likely to make mistakes when the context contains their errors from prior turns. Self-conditioning does not reduce by just scaling the model size. In contrast, recent thinking models do not self-condition, and can also execute much longer tasks in a single turn. We conclude by benchmarking frontier thinking models on the length of task they can execute in a single turn. Overall, by focusing on the ability to execute, we hope to reconcile debates on how LLMs can solve complex reasoning problems yet fail at simple tasks when made longer, and highlight the massive benefits of scaling model size and sequential test-time compute for long-horizon tasks.
[…]
@@ 1 Introduction
Is continued scaling of compute for Large Language Models (LLMs) economically justified given diminishing marginal gains? This question lies at the heart of the ongoing debate on the viability of continued massive investments in LLMs. While scaling laws show diminishing returns on metrics like test loss, the economic potential of LLMs might arise from automating long, multi-step tasks (METR, 2025). However, long-horizon tasks have been the Achilles’ heel of Deep Learning. We saw impressive self-driving demos take over a decade to translate to reliability in long-distance driving. Vision models can generate impressive images, and yet consistency over long videos remains an unsolved challenge. As the industry races to build agents that tackle entire projects, not just isolated questions, a fundamental question arises: How can we measure the number of steps an LLM can reliably execute?
Figure 1: A Summary of our contributions. We note that diminishing returns can enable exponential gains in horizon length (that is the length of tasks a model can complete). We design a simple task that isolates the capability of long-horizon execution in language models, ablating the need for knowledge and planning. We find that frontier models benefit considerably from both scaling model size and test-time compute when executing long horizon tasks.
LLM failures on simple, but long tasks have been considered a fundamental inability to reason (Mirzadeh et al., 2024). Despite massive improvements on complex reasoning benchmarks, Shojaee et al. (2025) claim thinking models (Guo et al., 2025) only give an “illusion of thinking”, as they eventually fail when the task is made longer. These results have sparked much debate in the community, which we think can be resolved by decoupling the need for planning and execution in reasoning or agentic tasks. Planning involves deciding what information to retrieve or tools to use and in which order, while execution involves carrying out the plan. In Shojaee et al. (2025), the LLM clearly knows the plan, as it initially follows it for many steps correctly. We posit that the eventual failures are in execution—as the task gets longer, the model is more likely to make a mistake in executing the plan. Although much attention has been paid to LLM planning abilities (Kambhampati et al., 2024), execution remains an understudied challenge, despite being increasingly important as LLMs begin to be used for long reasoning and agentic tasks.
@@
Does Scaling have Diminishing Returns? We observe that diminishing improvements in single-step accuracy can compound, leading to exponential growth in the length of task a model can complete. Traditionally, scaling model size is assumed to increase capacity to store parametric knowledge or search for plans. Yet, even when the required knowledge and plan are explicitly provided, empirically we find that scaling model size leads to large improvements in the number of turns a model can execute successfully.
The Self-Conditioning Effect. One might assume that failures on long tasks are simply due to the compounding of a small, constant per-step error rate. However, we find that the per-step error rate itself rises as the task progresses. This is in contrast to humans, who typically improve at executing a task with practice. We hypothesize that as a significant fraction of model training is to predict the most likely next token given its context, conditioning models on their own error-prone history increases the likelihood of future errors. We test this by controlling the error rate in the history shown to the model. As the error rate in the history is increased, we observe a sharp degradation in subsequent step accuracy, validating that models self-condition. We show self-conditioning leads to degradation in model performance in long-horizon tasks beyond previously identified long-context issues, and unlike the latter, is not mitigated by scaling model size.
The Impact of Thinking. We find recent thinking models are not affected by prior mistakes, fixing self-conditioning. Further, sequential test time compute greatly improves the length of task a model can complete in a single turn. Where without CoT, frontier LLMs like DeepSeek-V3 fail at performing even two steps of execution, its thinking version R1 can execute 200, highlighting the importance of reasoning before acting (Yao et al., 2023). We benchmark frontier thinking models, and find GPT-5 thinking (codename “Horizon”) can execute over 1000 steps, far ahead of the next best competitor, Claude-4-Sonnet at 432.
The “jagged frontier” (Dell’Acqua et al., 2023) of LLM capabilities remains fascinating yet confusing. Unlike traditional machines, LLMs are more susceptible to failure when used for executing repetitive tasks. Thus, we argue execution failures in long tasks should not be misinterpreted as the inability to reason or plan. We show long-horizon execution improves dramatically by scaling model size and sequential test time compute. If the length of tasks a model can complete indicates its economic value, continued investment in scaling compute might be worth the cost, even if short-task benchmarks give the illusion of slowing progress.
@@ 2 Formulation
In an agentic or reasoning task, the model begins in an initial state (based on the first input) and has to perform a sequence of steps to reach the final goal. A long-horizon task requires a large number of steps, where the task length is the number of steps needed to complete it. We define the following metrics to evaluate performance:
Step Accuracy. It measures the fraction of samples where the state update from step i1i-1 to step $i$ is correct, regardless of the correctness of the model’s state at step i1i-1.
Turn Accuracy. A turn is a single interaction with the model, which may require executing multiple steps. Turn Accuracy measures the fraction of samples where the state update from turn t1t-1 to turn $t$ is correct, regardless of the correctness of the model’s state at turn t1t-1.
Turn Complexity ($K$). It is defined as the number of steps the model has to execute per turn.
Task Accuracy. It measures the fraction of samples in which the model can complete a task of $i$ steps without making any mistakes in the process.
@@
Horizon Length ($H_{s}$). We define the horizon length of a model given a success rate threshold $0\leq s\leq 1$ as the first step $i$ where the model’s mean task accuracy across samples drops below $s$. It can be interpreted as: the model can perform a task of length $H_{s}$ without making mistakes, with probability $s$. We use $s=0.5$ unless otherwise specified, in analogy to Kwa et al. (2025).
Figure 2: Growth of Horizon Length. The length of task a model can perform at more than 50% accuracy grows faster than exponential as a function of the step accuracy after the 70% mark.
@@
2.1 Diminishing returns in Step Accuracy yield exponential gains on Horizon Length
@@
We begin by analyzing the relationship between a model’s single-step accuracy and its horizon length. To obtain a mathematical relation, we make two simplifying assumptions similar to LeCun (2023). First, we assume a model’s step accuracy remains constant over the task. Second, we assume a model does not self-correct, meaning any single error leads to task failure. We assume this only for the analysis here, which is illustrative and provides useful intuition. Our empirical analysis goes beyond this, investigating how LLMs, in fact, do not exhibit constant step accuracy for long horizon execution, and may correct mistakes.
@@
We plot this growth function in Figure˜2 for $s=0.5$. Note how after the step accuracy crosses 70%, small gains in step accuracy lead to faster than exponential improvement in horizon length. This derivation shows that even if accuracy improvements seem to slow down on question answering benchmarks, which typically have short tasks (METR, 2025), one can still mathematically expect large gains on longer tasks.
For example, on software engineering tasks, Kwa et al. (2025) empirically observed that the horizon length at $s=0.5$ of frontier models is growing exponentially, doubling every 7 months. Using our result above, in  Figure˜1 we show that such exponential growth in horizon length occurs even in a regime of diminishing returns on step accuracy. If we set $s=0.5$, we obtain $H_{0.5}=-\frac{\ln(2)}{\ln(p)}$. As such, the step-accuracy $p$ required to sustain exponential growth in $H_{0.5}$ over time ($x$) is $2^{\frac{-1}{2^{x}}}$, which is indeed a diminishing function.
We note that human labor is often compensated for its time. If the economic value of an agent also arises from the length of tasks it can complete, single-turn or short task benchmarks may be an illusory reference for evaluating the benefits of further investment in LLM compute. They might give a mirage of slowing progress, while the length of tasks a model can complete, which we think is a better indicator of economic value, continues to grow fast.
Figure 3: Overview of our framework. (Left) Our framework models long-horizon tasks as a sequence of retrieve-then-compose steps. (Right) We design a simple task that decouples planning from execution: in each turn, we provide the model the plan as key(s), asking it to retrieve their value(s), and compose them to maintain a running sum.
@@ 2.2 Isolating execution by decoupling planning and knowledge
We now describe how we measure long-horizon execution empirically. First, as a motivating example, consider an agent for the popular, economically valuable, task of booking flights. Upon receiving a search result, it must evaluate the surfaced flights to determine which one to book. The plan for assessing a single flight option may involve a sequence of actions, such as viewing detailed information, verifying that the flight timings, baggage allowance, and airline reviews align with user preferences, applying any available discounts or reward programs, and ultimately making a selection based on cost and travel time. Each of these individual steps requires retrieving some information, and composing it with the existing information state to eventually evaluate one flight option, and both of these operations require knowledge. The successful evaluation of multiple flight options constitutes the execution of this plan until a final booking decision is made.
In this work, we focus on execution, as we argue that it is a critical component of long-horizon capabilities. Execution has traditionally received less attention than capabilities such as reasoning, planning, and world knowledge, which have been the primary focus of LLM capability discussions. This relative neglect is significant to the extent that failures in execution have been misattributed to limitations in reasoning or planning capabilities (Shojaee et al., 2025; Khan et al., 2025). This perception may stem from the view that execution is a straightforward or mundane task. After all, this is what machines have been historically good at. Humans, too, are quite reliable at executing a task once they learn how to do it, even improving with practice. However, as LLMs do not come with correctness guarantees, we posit that execution can be surprisingly challenging for an LLM over a long horizon. We hypothesize that:
Even if reasoning, planning, and world knowledge are perfected,
LLMs will still make mistakes in execution over a long-horizon.
To demonstrate this, we isolate execution failures by explicitly providing the requisite knowledge and plan. We chain the retrieve-then-compose step motivated in the flight-selection agent example above. Each step involves retrieving relevant information or a tool specified by the plan and then composing its output to update the current state. The plan is deciding what to retrieve and how to compose it, whereas execution is actually performing those operations. This fits a natural abstraction—a key-value dictionary. The key serves as one step of a plan specifying what knowledge to retrieve, or tool to call, while the value represents the knowledge or tool output, which then has to be composed with the current state. In our study, we provide the plan as the keys in each query, eliminating the need for planning abilities from the LLM. We also provide the key-value dictionary in context, removing any dependency on the model’s parametric knowledge. With this design, we directly control two important axes that multiply to obtain the task length (number of retrieve-then-compose steps): the number of turns, and the turn complexity ($K$). The turn complexity can be varied by changing the number of keys queried per turn.
@@ 3 Experiments
Setup. As illustrated in Figure˜3, we provide the model with the needed knowledge, a fixed, in-context dictionary $\mathcal{D}:\mathcal{V}\to\mathbb{Z}$, where $\mathcal{V}$ is a vocabulary of common five-letter English words and values are integers sampled uniformly from [99,99][-99,99]. The initial state is $S_{0}=0$. In turn $t\in\{1,\dots,T\}$, the model receives an explicit plan $P_{t}=\{k_{t,1},\dots,k_{t,K}\}$, which is a set of $K$ keys sampled from $\mathcal{V}$. For each turn $t$, the model must execute this plan, which requires updating the state, $S_{t}$ to maintain a running sum of values for all past queried keys. This requires the retrieve-then-compose steps defined above:
@@
We choose short English words and two-digit integers to minimize errors arising from tokenization. The state transition here is Markovian, depending only on $S_{t-1}$ and $P_{t}$. The task is extremely simple, by design, to isolate long-horizon execution by minimizing the knowledge needed for the retrieval and composition operation. More details, including the exact prompt, are provided in Appendix˜E. We analyze empirical performance on the individual retrieval, and composition operations in Appendix˜D and the format following errors in Appendix˜F.
Figure 4: Scaling model size non-diminishingly improves the number of turns it can execute. We vary the model size and study both full task (a) and turn-wise accuracy (b) as the number of turns increases. Bold lines are a running average of accuracy over 5 turns. The dotted lines (turn-wise accuracy) in (b) show single-step accuracy for our task is 100% for all except the smallest models. Yet, as the number of turns increases, the performance gap between small and large models widens (a), with the latter having significantly more horizon length (c).

### v2

@@ Abstract
Does continued scaling of large language models (LLMs) yield diminishing returns? In this work, we show that short-task benchmarks may give an illusion of slowing progress, as even marginal gains in single-step accuracy can compound into exponential improvements in the length of tasks a model can successfully complete. Then, we argue that failures of LLMs when simple tasks are made longer arise from mistakes in execution, rather than an inability to reason. So, we propose isolating execution capability, by explicitly providing the knowledge and plan needed to solve a long-horizon task. First, we find that larger models can correctly execute significantly more turns even when small models have near-perfect single-turn accuracy. We then observe that the per-step accuracy of models degrades as the number of steps increases. This is not just due to long-context limitations—curiously, we observe a self-conditioning effect—models become more likely to make mistakes when the context contains their errors from prior turns. Self-conditioning does not reduce by just scaling the model size. But, we find that thinking mitigates self-conditioning, and also enables execution of much longer tasks in a single turn. We conclude by benchmarking frontier thinking models on the length of tasks they can execute in a single turn. Overall, by focusing on the ability to execute, we hope to reconcile debates on how LLMs can solve complex reasoning problems yet fail at simple tasks when made longer, and highlight the massive benefits of scaling model size and sequential test-time compute for long-horizon tasks.
Code           ?  Dataset
@@ 1 Introduction
Is continued scaling of compute for Large Language Models (LLMs) economically justified given diminishing marginal gains? This question lies at the heart of the ongoing debate on the viability of continued massive investments in LLMs. While scaling laws show diminishing returns on metrics like test loss, the true economic potential of LLMs might arise from automating long, multi-step tasks (METR, 2025). However, long-horizon tasks have been the Achilles’ heel of Deep Learning. We see recent vision models generating impressive images, yet consistency over long videos remains an unsolved challenge. As the industry races to build agents that tackle entire projects, not just isolated questions, a fundamental question arises:
How can we measure the number of steps an LLM can reliably execute?
Figure 1: A summary of our contributions. Our work measures long-horizon execution, finding large benefits from scaling model size and sequential test-time compute. We identify a failure mode where models self-condition on their own errors, degrading future performance.
LLM failures on simple, but long tasks have been considered a fundamental inability to reason (Mirzadeh et al., 2024). Despite massive improvements on complex reasoning benchmarks, Shojaee et al. (2025) claim thinking models (Guo et al., 2025) only give an “illusion of thinking”, as they eventually fail when the task is made longer. These results have sparked much debate in the community, which we think can be resolved by decoupling the need for planning and execution in reasoning or agentic tasks. Planning involves deciding what information to retrieve or tools to use and in which order, while execution involves carrying out the plan. In Shojaee et al. (2025) setup, the LLMs know the correct plan, as they initially follow it correctly for many steps. We posit that the eventual failures are in execution—as the task gets longer, the model is more likely to make a mistake in executing the plan. Although much attention has been paid to LLM planning abilities (Kambhampati et al., 2024), execution remains an understudied challenge, despite being increasingly important as LLMs begin to be used for long reasoning and agentic tasks.
@@
Does Scaling have Diminishing Returns? We observe that diminishing improvements in single-step accuracy can compound, leading to exponential growth in the length of task a model can complete. Traditionally, scaling model size is assumed to increase capacity to store parametric knowledge or search for plans. Yet, even when the required knowledge and plan are explicitly provided, we find that scaling model size leads to large improvements in the number of turns a model can execute successfully.
The Self-Conditioning Effect. One might assume that failures on long tasks are simply due to the compounding of a small, constant per-step error rate and context length issues. However, we find that the per-step error rate itself rises as the task progresses. This is in contrast to humans, who typically improve at executing a task with practice. We hypothesize that, as a significant fraction of model training is to predict the most likely next token given its context, conditioning models on their own error-prone history increases the likelihood of future errors. We test this by controlling the error rate in the history provided to the model. As the error rate in the history is increased, we observe a sharp degradation in subsequent step accuracy, validating that models self-condition. We show how self-conditioning leads to degradation in model performance in long-horizon tasks beyond previously identified long-context issues, and unlike the latter, is not mitigated by scaling model size.
The Impact of Thinking. We find that recent thinking models are not affected by prior mistakes, fixing the self-conditioning effect. Further, sequential test time compute also significantly improves the length of task a model can complete in a single turn. Where, without chain-of-thought prompting, even frontier LLMs like DeepSeek-V3 fail at performing even four steps of execution, and its thinking version R1 can execute over 100 steps, highlighting the importance of reasoning before acting (Yao et al., 2023). We benchmark frontier thinking models, and find GPT-5 thinking (codenamed “Horizon”) can execute over 2100 steps, far ahead of the next best competitor, Claude-4 Sonnet at 432.
The “jagged frontier” (Dell’Acqua et al., 2023) of LLM capabilities remains fascinating yet confusing. Unlike traditional machines, LLMs are more susceptible to failure when used for executing repetitive tasks. Thus, we argue that execution failures in long tasks should not be misinterpreted as the inability to reason. We show that long-horizon execution improves dramatically by scaling model size and sequential test time compute. If the length of tasks a model can complete indicates its economic value, continued investment in scaling compute might be worth the cost, even if short-task benchmarks give the illusion of slowing progress.
@@ 2 Formulation
First, we define key capabilities involved in an agentic or reasoning task. As a motivating example, consider an agent for the task of booking flights. Upon receiving a search result, the agent must reason to choose a flight that aligns with user preferences. One plan for this reasoning task could be:
For each flight, verify the flight timings, baggage allowance, and airline reviews. Then apply any available discounts or reward programs, and finally select a flight based on cost and travel time.
Each of these individual steps requires two operations: retrieving some information, and composing it with the existing information state, until the goal of choosing the final flight is reached. Both these operations require knowledge, potentially tacit, about how to perform them. Execution is carrying out this plan, a sequence of retrieve-then-compose steps, until a final booking is made. We formalize these terms for our work as follows:
Key Terms Planning. Deciding what steps to take, and in what sequence. Execution. Carrying out the steps decided in the plan. Knowledge. Information about different types of steps, and how to compose them. A reasoning task. Requires planning the steps needed to solve it, and then executing them. An agentic task. Requires planning what actions to take, and then executing them.
In this work, we focus on execution, as we argue that it is a critical component of long-horizon capabilities. Execution has traditionally received less attention (Stechly et al., 2024) than capabilities such as reasoning, planning, and world knowledge, which have been the primary focus of LLM capability discussions. In fact, failures in execution have been misattributed to limitations in reasoning or planning capabilities (Shojaee et al., 2025; Khan et al., 2025). This perception may stem from the view that execution is straightforward or mundane. For example, once we as humans learn how to do a task, we are quite reliable at executing it, even improving with practice. However, as LLMs do not come with correctness guarantees, we posit that just execution over a long horizon can surprisingly be a challenge. We hypothesize that:
Even if planning and world knowledge are perfected,
LLMs will still make mistakes in execution over a long-horizon.
@@
In an agentic or reasoning task, the model begins in an initial state (based on the first input) and has to perform a sequence of steps to reach the final goal. A long-horizon task requires a large number of steps, with the task length being the number of steps needed to complete it. We define the following metrics to evaluate performance:
@@
Evaluation Metrics Step Accuracy. Measures the fraction of samples where the state update from step i1i-1 to step $i$ is correct, regardless of the correctness of the model’s state at step i1i-1. Turn Accuracy. A turn is a single interaction with the model, which may require executing multiple steps. Turn Accuracy measures the fraction of samples where the state update from turn t1t-1 to turn $t$ is correct, regardless of the correctness of the model’s state at turn t1t-1. Turn Complexity ($K$). Defined as the number of steps the model has to execute per turn. Task Accuracy. Measures the fraction of samples in which the model can complete a task of $i$ steps without making any mistakes in the process. Horizon Length ($H_{s}$). Given a success rate threshold $0\leq s\leq 1$, the horizon length is the first step $i$ where the model’s mean task accuracy across samples drops below $s$. It can be interpreted as: the model can perform a task of length $H_{s}$ without making mistakes, with probability $s$. We use $s=0.5$ unless otherwise specified, like Kwa et al. (2025).
@@
2.1 Diminishing Returns in Step Accuracy Compound Over a Long Horizon
@@
We begin by analyzing the relationship between a model’s single-step accuracy and its horizon length. Note that this analysis applies not just to execution, but rather to any general long-horizon task. To obtain a mathematical relation, we make two simplifying assumptions similar to LeCun (2023). First, we assume a model’s step accuracy remains constant over the task. Second, we assume a model does not self-correct, meaning any single error leads to task failure. We assume this only for the analysis here, which is illustrative and provides useful intuition. Our subsequent empirical analysis goes beyond this, investigating how LLMs, in fact, do not exhibit constant step accuracy for long-horizon execution, and may also correct their mistakes.
Figure 2: Growth of Horizon Length. The length of task a model can perform grows hyperbolically in the high accuracy regime.
@@
This shows that the horizon length grows hyperbolically with the step accuracy. We illustrate this growth in Figure˜2 across different values for the success rate $s$. Notice the sharp growth in horizon length beyond 80% single-step accuracy, performance that frontier models now achieve on many question-answering benchmarks (Vendrow et al., 2025), which can be considered short tasks.
We note that human labor is often compensated for its time. If the economic value of an agent also arises from the length of tasks it can complete, single-turn or short task benchmarks may be misleading for evaluating the benefits of further investment in LLM compute. While these benchmarks reveal genuine diminishing returns at the step level, they understate the compounding benefits that emerge over long-horizons. Beyond a threshold, small improvements in step accuracy can translate into success at rapidly increasing task lengths, which may provide a more faithful indicator of economic value.
For example, in METR’s horizon length plot on software engineering tasks (Kwa et al., 2025), it was empirically observed that the horizon length at $s=0.5$ of frontier models is growing exponentially, doubling every 7 months. Using our result above, in  Figure˜1 we show that such exponential growth in horizon length occurs even in a regime of diminishing returns on step accuracy. If we set $s=0.5$, we obtain $H_{0.5}=-\frac{\ln(2)}{\ln(p)}$. As such, the step-accuracy $p$ required to sustain exponential growth in $H_{0.5}$ over time ($t$) is $2^{\frac{-1}{2^{t}}}$, which is indeed a diminishing function.
@@ 2.2 Isolating execution by decoupling planning and knowledge
We now describe how we measure long-horizon execution empirically. We isolate execution failures by explicitly providing the requisite knowledge and plan. We study the chaining of the retrieve-then-compose step motivated in the flight-selection agent example earlier. Each step involves retrieving relevant information or a tool specified by the plan and then composing its output to update the current state. The plan is deciding what to retrieve and how to compose it, whereas execution is actually performing those operations. This fits a natural abstraction—a key-value dictionary. The key serves as one step of a plan specifying what knowledge to retrieve, or tool to call, while the value represents the knowledge or tool output, which then has to be composed with the current state. In our study, we provide the plan as the keys in each query, eliminating the need for planning abilities from the LLM. We also provide the key-value dictionary in context, removing any dependency on the model’s parametric knowledge. With this design, we directly control two important axes that multiply to obtain the task length (number of retrieve-then-compose steps): the number of turns, and the turn complexity ($K$). The turn complexity can be varied by changing the number of keys queried per turn.
Figure 3: Overview of our framework. (Left) Our framework models long-horizon tasks as a sequence of retrieve-then-compose steps. (Right) We design a simple task that decouples planning from execution: in each turn, we provide the model the plan as key(s), asking it to retrieve their value(s), and compose them to maintain a running sum. We control the number of turns and turn complexity (keys per query).
@@ 3 Experiments
We design a simple task where even language models with 4 billion parameters can achieve high accuracy, to isolate the capability of long-horizon execution.
Setup. As illustrated in Figure˜3, we provide the model with the needed knowledge, a fixed, in-context dictionary $\mathcal{D}:\mathcal{V}\to\mathbb{Z}$, where $\mathcal{V}$ is a vocabulary of common five-letter English words and values are integers sampled uniformly from [99,99][-99,99]. The initial state is $S_{0}=0$. In turn $t\in\{1,\dots,T\}$, the model receives an explicit plan $P_{t}=\{k_{t,1},\dots,k_{t,K}\}$, which is a set of $K$ keys sampled from $\mathcal{V}$. For each turn $t$, the model must execute this plan, which requires updating the state, $S_{t}$, to maintain a running sum of values for all past queried keys. This requires the retrieve-then-compose steps:
@@
We choose short English words and two-digit integers to minimize errors arising from tokenization. We provide few-shot examples to clarify the task. More details, including the exact prompt, are provided in Appendix˜E. We also disentangle performance on the individual retrieval and composition operations, finding that models have much higher accuracies on each of them alone (Appendix˜D).